OpenStudy (anonymous):

i need to Find the value of 5u+5 given that 2u-7=3 . Simplify your answer as much as possible.

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

first we solve the first equation 2u-7=3. We add seven to both sides to get 2u=10, and we get u by itself by dividing both sides by 2, u=5. Now that we know that u=5, we can plug it in to the other equation. 5(5)+5=?. We multiply the 5 on the outside of the parentheses by the 5 on the inside, and we get 25+5=?. 25+5=30, so 30 is our final answer.

OpenStudy (anonymous):

okay